OpenStudy (anonymous):

find the area of this region using integrals. y=x and y=x^2

OpenStudy (anonymous):

draw a picture first http://www.wolframalpha.com/input/?i=y%3Dx^2%2C+y%3Dx

OpenStudy (loser66):

@satellite73 quickly without any explanation, ha!!

OpenStudy (anonymous):

y=x to find the point of intersection,put the value of y in y=x^2 x^2=x x^2-x=0 x(x-1)=0 x=0,1  \[Required area=\int\limits_{0}^{1} x dx-\int\limits_{0}^{1} x ^{2}dx\] \[(area=\int\limits_{0}^{1}y dx)\] calculate.